OpenStudy (anonymous):

x-2=-2 -x+2y=1 solve by elimination

OpenStudy (ashes.boom):

from what you have written it looks like it would be y=1/2. Heres how i got it x - 2 = -2 -x + 2y = 1 the x variable cancels out 0 + 2y - 2 =(-2+ 1) 2y - 2 = -1 add 2 to both sides +2 +2 2y = 1 divide by 2 to get the y variable /2 /2 y = 1/2

OpenStudy (ashes.boom):

Does that help at all